# Effect of invisible neutrino decay on the neutrino oscillation at long-baselines

### v1

@@ 1 Introduction
In contrast, processes involving neutrino decay in physics beyond the Standard Model (BSM) such as $\nu_{i}\rightarrow\nu_{f}+X_{f}$ are much less constrained. Here $\nu_{i}$ denotes the mass states (i=1,2,3), and $\nu_{f}$ and $X_{f}$ are the final state particles. From a phenomenological perspective, there are two types of neutrino decay: visible decay and invisible decay. In visible decay, the neutrinos in the final state are active and can be detected. This type of decay would result in the observable event rates at the detector being impacted by the decay products, which would have lower energies than the parent particle. A typical signature of visible neutrino decay would be a concentration of events at low energiesKim:1990km ; Acker:1992eh ; Lindner:2001fx ; Coloma:2017zpg ; Gago:2017zzy . In invisible decay, the neutrino in the final state is not observable in the detector. This could occur if the neutrino in the final state is a light sterile neutrino (lighter than the active ones, allowing for the decay kinematically). Another possibility is that the decay product involves an active neutrino, which, due to its low energy, is unobservable.
We are interested in discussing the phenomenon of invisible neutrino decay. In this particular scenario, a beam of relativistic neutrinos ($\nu_{i}$) with a lifetime of $\tau_{i}$ and mass $m_{i}$ is subject to depletion as a consequence of invisible decay by a factor of
@@
The decay of ultrahigh energy astrophysical neutrinos has been considered by many authorsBeacom:2002vi ; Maltoni:2008jr ; Pakvasa:2012db . A recent study performed in Pagliaroli:2015rca states that the IceCube experiment can reach a sensitivity of $\tau/m>10$ s/eV for both hierarchies for 100 TeV neutrinos coming from a source at a distance of 1 Gpc. The implications of invisible neutrino decay for future experiments like JUNOHe:2014zwa , RENO-50Kim:2014rfa , DUNE DUNE:2015lol , T2HK, and T2HKK Hyper-Kamiokande:2016dsw ; Hyper-Kamiokande:2016srs have been investigated.
@@ 5.4 Sensitivity in $\theta_{23}-\delta_{CP}$ Plane
  • The allowed contours are larger in size for P2O, signifying a lower sensitivity in $\theta_{23}$ and $\delta_{CP}$ than the DUNE-like setup for both decay and no-decay cases.
@@
Figure 12: Sensitivity in the test $\theta_{23}-\delta_{CP}$ plane for true values of $\theta_{23}=41^{\circ},49^{\circ}$, and $\delta_{CP}=-90^{\circ},0^{\circ},90^{\circ}$ for both normal and inverse hierarchy. The shaded contours are for the decay case, and the contours encompassed by curves are for the no-decay case. The dark green and light green correspond to $3\sigma,5\sigma$ contours in the combined analysis of P2O+LArTPC .
@@ Acknowledgements.
AC acknowledges the Ramanujan Fellowship (RJF/2021/000157), of the Science and Engineering Research Board of the Department of Science and Technology, Government of India and CERN EP-Nu group. SG acknowledges the J.C. Bose Fellowship (JCB/2020/000011) of the Science and Engineering Research Board of the Department of Science and Technology, Government of India. It is to be noted that this work has been done solely by the authors and is not representative of the DUNE collaboration.

### v2

# Effect of invisible neutrino decay on neutrino oscillation at long baselines

@@ 1 Introduction
In contrast, processes involving neutrino decay in physics beyond the Standard Model (BSM) such as $\nu_{i}\rightarrow\nu_{f}+X_{f}$ are much less constrained. Here $\nu_{i}$ denotes the mass states (i=1,2,3), and $\nu_{f}$ and $X_{f}$ are the final state particles. From a phenomenological perspective, there are two types of neutrino decay: visible decay and invisible decay. In visible decay, the neutrinos in the final state are active and can be detected. This type of decay would result in the observable event rates at the detector being impacted by the decay products, which would have lower energies than the parent particle. A typical signature of visible neutrino decay would be a concentration of events at low energiesKim:1990km ; Acker:1992eh ; Lindner:2001fx ; Coloma:2017zpg ; Gago:2017zzy . In invisible decay, the neutrino in the final state is not observable in the detector. This could occur if the neutrino in the final state is a light sterile neutrino (lighter than the active ones, allowing for the decay kinematically). Another possibility is that the decay product involves an active neutrino, which, due to its low energy, is unobservable. We are interested in discussing the phenomenon of invisible neutrino decay. In this particular scenario, a beam of relativistic neutrinos ($\nu_{i}$) with a lifetime of $\tau_{i}$ and mass $m_{i}$ is subject to depletion as a consequence of invisible decay by a factor of
@@
[…]
@@ 5.4 Sensitivity in $\theta_{23}-\delta_{CP}$ Plane
  • The allowed contours are larger for P2O, signifying a lower sensitivity in $\theta_{23}$ and $\delta_{CP}$ than the DUNE-like setup for both decay and no-decay cases.
@@
Figure 12: Sensitivity in the test $\theta_{23}-\delta_{CP}$ plane for true values of $\theta_{23}=41^{\circ},49^{\circ}$, and $\delta_{CP}=-90^{\circ},0^{\circ},90^{\circ}$ for both normal and inverse hierarchy. The shaded contours are for the decay case, and the contours encompassed by curves are for the no-decay case. The dark green and light green correspond to $3\sigma,5\sigma$ contours in the combined analysis of P2O+LArTPC setups.
@@ Acknowledgements.
AC acknowledges the Ramanujan Fellowship (RJF/2021/000157), of the Science and Engineering Research Board of the Department of Science and Technology, Government of India, and CERN EP-Nu group. SG acknowledges the J.C. Bose Fellowship (JCB/2020/000011) of the Science and Engineering Research Board of the Department of Science and Technology, Government of India. It is to be noted that this work has been done solely by the authors and is not representative of the DUNE collaboration. SP appreciates the guidance of Dr. Monojit Ghosh regarding the writing of the glb file for the P2O experiment.